I dont really understand this qeustion so if someone would able to help me it would be very much appreicated.
Explain in details and in your own words why the two equations below are incorrect. Then show the correct algebra, working from the LHS.
(i) 2/x + 1/2y = 3/x+2y

(ii) (x – 2)^3 = x^3 – 2^3

Answer by stanbon(75887)

(i) 2/x + 1/2y = 3/x+2y
-----------
Adding the numerators and Adding the denominators
is not the way to add fractions.
You need a common denominator to add fractions.
-----
Correct procedure:
2/x + 1/(2y)
least common denominator is 2xy
---
Rewrite each fraction with the lcd as its denominator:
(2(2y))/lcd + (1*x)/lcd
Combine the numerators over the lcd:
= (4y+x)/(2xy)
=======================================
(ii) (x – 2)^3 = x^3 – 2^3
Raising each term to the 3rd power is not the
way to raise a binomial to a power.
---
(x-2)^3 = x^3 + 3x^2(-2) + 3x(-2)^2 + (-3)^3
--------
= x^3 -6x^2 +12x - 27
============================
Cheers,
Stan H.
==========================